by TEI TEI

02/27/06 – JCI reported today that all final documents are on their way to Tony and they will train the Admin staff on outside access for this project and take care of small punchlist items.

02/23/06 – Final paperwork received. No significant punchlist items other than to answer a couple of questions for us on the control system parameters on the job. TTC will closeout the project. TEI UPS final payapp, closeout paperwork and O&M’s to Tony Matre today. Final Punchlist items included: Item #1: What controls the EF-1 operation? Is it just the rooftop units? That appears to be the only thing telling the fan when to run. The problem is that aside from the OA enable setpoint that EF status should be tied to all the air handlers for relief during economizer mode not just the RTU’s they are almost incidental to its operation. Please clarify. Item #2: The speed drive has been hunting continuously since Mike Asbrock made the latest programming changes. Please investigate and remedy. Item #3: Are you providing a laptop on this project for Joe Merkt? Regardless, we need training to show Joe how to pull the ADS server up from outside the office. Have you coordinated with HamCo IT for internet access or port security control into the building? Does Joe and his staff understand how to view from a remote location?

01/18/06 – Wayne Jackson reported the JCI has been on site working on the punchlist this week. Joe Merkt confirmed they had indeed been at the admin building. The licensed admin software is not on site but JCI will find it and deliver it to Joe. A full backup of the JCI system was completed today. TTC will schedule final punchlist visit for next week to wrap up the job. JCI will provide final documents and billing by end of January.

01/10/06 – TTC emailed JCI to inquire about taking care of punchlist emailed on 12/13/05.

12/13/05 – TTC emailed a partial punchlist to JCI. No response has been received or questions answered yet. AH-6 has been repaired by JCI at this time.

12/06/05 – TTC stopped by the building and began the punchlist today. It is not complete and a second trip will be needed. TTC to schedule in the near future.

12/02/05 – Admin Punchlist has been scheduled for 12/06/05. TTC will distribute soon thereafter.

10/24/05 – JCI is complete and TTC is supposed to contact Joe and setup final punchlist for the project.

10/04/05 – JCI was supposed to complete the job on Friday. TTC will check and find out and will advise the County.

10/03/05 – JCI did not get finished with the DDC project as promised. They believe they will be done this Friday with graphics and 10th floor upgrades 10/07/05. TTC has to follow up with the punchlist as soon as they are complete.

09/19/05 – JCI emailed back and confirmed that the deadline is 10/01/05 and they will be done. TTC told Tony that TE will punchlist the job and get punchlist complete by mid-October.

09/16/05 – TTC emailed JCI about working onsite and meeting the 10/01 deadline.

09/08/05 – JCI emailed to commit to an 10/01/05 completion date on all remaining items on the DDC project. Parts should be in late this week and they can start this week or next. They will contact Joe for on-site work.

09/01/05 – TTC emailed JCI today about final completion again.

08/31/05 – Joe called about getting a final date for completion. This project has gone on too long and Tony wants a completion date that includes the change order cost.

08/24/05 – TTC contacted JCI with all the change order and allowance usage information to get final approval for doing that work. Once complete no allowance money will be left and JCI should bill Joe directly for remaining cost. TE will advise and information comes back.

08/23/05 – JCI and Wayne Jackson have sent in a revised current sensor change order without all low voltage wiring ran in control. The new cost is only $5700. TTC passed along to County for approval. TTC had a voicemail from Mike Asbrock on Monday stating that the hot water pumps were alarmed but not the boilers per say and the only way to know the heating boiler was in alarm was by a decrease in hot water temperature.

08/08/05 – JCI called in response to my inquiry about project status and provided the following report. The project is still about one month away from being complete. Each floor has a reset schedule that they were unaware of that appears to have been added after the original programming and they have to go back and put that into each DX unit. The will have this done in a couple of weeks – their technicians is on vacation this week. The graphics aren’t started yet but they will have their technician do that in a couple of days – but they have to completely wipe the 630 main program off and start over to make it right. They are working on re-pricing the current sensors because their electricians are 100% complete at the building and to bring them back that will cost mobilization, but I told them we still need a price and to coordinate with Joe the scope of what he wanted on the closed feedback controls. Wayne Jackson went on to say that they know they still have a lot of small cleanup work to complete and I told them that and to update the as-builts with all the new sequences.

08/03/05 – Joe called today and asked about revisiting the circuit sensor change order since the chiller job came in under budget. TTC sent JCI and email about final status and about what the cost would be to upgrade the controls and sensors to monitor actual running status of the equipment at the Admin. The old control system had open feedback also – so it didn’t get upgraded during the project. TTC also inquired about the graphics upgrade and when the project would be complete.

06/28/05 – TE received invoice # 01 from JCI but as Pat pointed out, it is not on an AIA form and it doesn’t include the correct paperwork for approval. TTC will contact JCI.

06/15/05 – Joe called today to report that Tony will not be accepting the current sensor change order. JM also reported that JCI was finished with floor 1-8 on all HVAC units. No graphics have been installed to date. TTC said that made sense because that will be the last part that Ed Braemer puts it when he arrives on site for commissioning and downloading.

06/09/05 – TE received update submittals showing the chilled water system for approval.

05/19/05 – JCI is onsite with ACPI on the 8th floor working on the FCU panel. A new DX9100 is being installed and ancillary equipment is being installed before the changeover takes place.

04/22/05 – TE has reviewed submittals and can approve after JCI answers a couple of small questions about AHU-1 and 3. Should leave TE on Monday.

04/06/05 – TE received submittals on controls from JCI. TE will begin review of entire packet for use on the new job. TE is still awaiting direction from the County on whether to accept the CO request for fan status on each air handler.

03/24/05 – Submittals to TE at end of next week. Electrician (ACPI) has pulled the N2 bus around the building. JCI thinks that they will be on site a couple of weeks after that. JCI is still awaiting an answer on circuit sensors.

03/03/05 – JCI turned in a $$9226 Change Order request to add current sensors to all the air handler fans. Apparently the existing system only has open loop control of these fans meaning that if the fans are commanded on, the system shows them as running, if they fail or don’t turn on, the DDC will not show them in alarm. Please advise on how to proceed on this change order request.

01/19/05 – TTC, Ed Braemer and Joe Merkt discussed additional air handler today but couldn’t resolve the numbering issue. We will meet onsite on Jan 24 to resolve. Joe believes that the quote includes all air handlers for this project.

01/12/05 – Joe Shutte called to tell us that an additional air handler on the first floor was discovered during field work that does not appear on any JCI control dwgs or as-builts, therefore it was not included in the state term price given to us for this project. TE informed them to match the existing sequence and give us a price to upgrade it on this job pending a decision from Facilities. We are awaiting the allowance/change order price.

11/05/04 – Pre Construction Meeting

09/20/04 – Tony and Diana requested an SOSS for this project today. TE completed the state term contract for JCI and has already sent it down.